A random sample of ten metal rods produced by a machine is taken. Each rod is tested for
hardness. The results, in suitable units, are as follows
525 520 522 524 518 520 519 525 527 516
Calculate an unbiased estimate of the population variance.
Find a 98% confidence interval for the population mean.
1
Expert's answer
2020-04-13T13:21:06-0400

n=10, x ={525, 520, 522, 524, 518, 520, 519, 525, 527, 516 }

"\bar{x}=\frac{\displaystyle\sum_{i=1}^n x_i}{n}" = (525+520+522+524+518+520+519+525+527+516)/10 = 521.6  

An unbiased estimate of the population variance.  

"\sigma^{2}=\frac{\displaystyle\sum_{i=1}^n (x_i -\bar{x})^{2}}{n-1} =\\= ((525-521.6)^2+(520-521.6)^2+(522-521.6)^2+...+(516-521.6)^2)/(10-1) = 12.71"

A 98% confidence interval for the population mean:

"\bar{x}-\Delta_x \le\bar{x}\le\bar{x}+\Delta_x"

"\u2206 _x= t *l,"

Find t in the t-Table : t(9,0.02) = 2.82,

"l=\sqrt{\smash[b]{\sigma^{2}/n}} =\sqrt{ 12.71/10}=" 1.1274,

"\Delta_x" = 1.1274*2.82 = 3.1793 ,

521.6- 3.1793 "\le\bar{x}\le" 521.6 +3.1793

518.42 "\le\bar{x}\le" 524.78